OpenStudy (df001):

Help Trig SinΘCosΘ+SinΘ/CosΘ+Cos^2Θ=TanΘ

OpenStudy (df001):

\[\frac{ Sin \theta Cos \theta + Sin \theta }{ Cos \theta + Cos ^{2} \theta }\]

OpenStudy (df001):

= Tanx

OpenStudy (df001):

pls help me

OpenStudy (anonymous):

factoring would help

OpenStudy (df001):

where can i begin

OpenStudy (anonymous):

then cancel

OpenStudy (mathmale):

The second term in the original equation is sin theta / cos theta.

OpenStudy (anonymous):

\[\frac{ab+b}{a+a^2}=\frac{b(1+a)}{a(1+a)}\] it is 100% algebra 0% trig

OpenStudy (mathmale):

Factor the numerator (in which both terms have sin x); Factor the denominator (in which both terms have cos x) See if that helps.

OpenStudy (mathmale):

Satellite73 and I are discussing the same approach (factoring), using different examples.

OpenStudy (df001):